OpenStudy (anonymous):

HELPP!!! will give medal :) Given parallelogram EFGH with perimeter 58, solve for EF EG and EH

OpenStudy (anonymous):



OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

I kno I have to find x but idk how I though I had the right idea but it was wrong

OpenStudy (raden):

x+x+x+3+x+3 = 58 combine the similar terms, then solve for x

OpenStudy (anonymous):

4x+6=58 4x=52 x=13

OpenStudy (raden):

yes, you are right!

OpenStudy (anonymous):

Ok thanks I think I got it from here :)

OpenStudy (raden):

so, EF = x = 13 EG = 2x = 2 *13 = 26 EH = x+3 = 13+3 = 16 wlcm